OpenStudy (anonymous):

Evaluate. cube root of 7 multiplied by the square root of 7 over sixth root of 7 to the power of 5 A.) 7-1 B.) 1 C.) 7 D.) 7 to the power of 5 by 3

OpenStudy (agreene):

it might be easier to convert all those roots to fractions and then use the identity \(a^xa^y=a^{x+y}\)

OpenStudy (anonymous):

\[\large \frac{\sqrt[3]{7} \times \sqrt{7}}{(\sqrt[6]{7})^5}\]

OpenStudy (midhun.madhu1987):